# supplement

### v1

@@ 1 Introduction
Nevertheless, synthesizing information from multiple sources is highly challenging. First, the prospective between-study heterogeneity complicates data integration. Such heterogeneity may manifest in various forms, including variations in genotype-phenotype associations (Frazier-Wood et al.,, 2013) and the differences in genetic variants across diverse ethnic populations (Gurdasani et al.,, 2019). Additionally, variability arising from inconsistent data collection and measurement across institutions further contributes to data heterogeneity (Kohane et al.,, 2021). Second, due to privacy or communication concerns, individual raw data often cannot be shared across different locations. Instead, only summary or intermediate statistics can be communicated among different sources. This constraint on data sharing may negatively impact the learning accuracy, exacerbating the challenges of data integration.
@@ 1.1 Related Works
In recent decades, significant progress has been made in the field of high-dimensional precision matrix estimation under sparsity, including penalized maximum likelihood procedures (Yuan and Lin,, 2007; Rothman et al.,, 2008; Lam and Fan,, 2009; Ravikumar et al.,, 2011), penalized linear regression approaches (Meinshausen and Bühlmann,, 2006; Sun and Zhang,, 2013; Ren et al.,, 2015), and regularized estimating equation methods (Yuan,, 2010; Cai et al.,, 2011; Cai et al., 2016b, ). The non-negligible bias induced by the regularization in these methods has prompted numerous debiasing strategies aimed at statistical inference on the entries of precision matrices (Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018; Chang et al.,, 2018).
@@
In a similar context to joint estimation, there is a substantial body of literature focusing on distributed analysis under distributional homogeneity. Examples include the one-shot averaging methods (Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Fan et al.,, 2019; Lv and Lian,, 2022), and the surrogate loss function methods (Wang et al.,, 2017; Jordan et al.,, 2019) with applications in distributed optimization (Fan et al.,, 2023), simultaneous inference (Yu et al.,, 2022), and non-smooth analysis (Tan et al.,, 2022). These methods achieve communication-efficiency in the sense that the dimensions of the transmitted statistics are proportional to those of target parameters. In addition, Chen et al., (2019) smooths the non-smooth objective and solves it by transmitting statistics with dimensions that are quadratic in the parameter dimension; see also Wang et al., (2019); Chen et al., (2024) for further applications. On the other hand, when model distributions are heterogeneous while target parameters remain homogeneous across sites, distributed analysis can be achieved by averaging (Zhao et al.,, 2016), weighted averaging (Gu and Chen,, 2023), or surrogate estimating equations (Duan et al.,, 2022).
@@ 1.2 Contributions
The newly established methodologies and theories in this article make several valuable contributions to both distributed statistical learning and integrative precision matrix analysis. First, the proposed HEAT and IteHEAT methods introduce novel methodological advances in distributed algorithms under heterogeneity. In comparison to existing communication-efficient algorithms developed under homogeneity, our approaches demonstrate flexibility and applicability across distributional heterogeneity scenarios. Additionally, our methods provide communication advantages over distributed algorithms designed for heterogeneous settings, such as Cai et al., (2022). Second, the proposed IteHEAT algorithm provides insights into the statistical refinement of higher-order errors and demonstrates potential for improving estimation accuracy in integrative analysis. Third, the proposed methods complement the field of multi-task learning for precision matrices. Unlike the typical joint estimation approaches that rely on solving highly non-smooth optimizations, HEAT and IteHEAT are optimization-free for integration, ensuring computational efficiency. Fourth, we establish the statistical optimality theories of HEAT and IteHEAT by deriving the rates of convergence and the corresponding minimax lower bounds. To the best of our knowledge, the current minimax theories are the first among the extensive literature on joint precision matrix estimation. The introduced lower bound techniques and constructions of least favorable matrices are of independent interest and make valuable contributions to the field.
@@ 2.1 Notations
For an integer $k>0$, let $[k]=\{1,2,\ldots,k\}$; denote by $0_{k}$ and $1_{k}$ the vectors of zeros and ones with length $k$, respectively. For a set $\mathcal{A}$, define $|\mathcal{A}|$ to be its cardinality. For a vector $a\in\mathbb{R}^{n}$ and a subset $\mathcal{A}\subset[n]$, denote by $a_{\mathcal{A}}=\left(a_{i}\right)_{i\in\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|}$ a subvector of $a$. Let $\mathcal{A}^{c}=[n]\setminus\mathcal{A}$, and define $a_{-\mathcal{A}}=a_{\mathcal{A}^{c}}\in\mathbb{R}^{n-|\mathcal{A}|}$, and $a_{-j}=a_{-\{j\}}$. For a matrix $A=\left(A_{i,j}\right)_{i\in[n_{1}],j\in[n_{2}]}\in\mathbb{R}^{n_{1}\times n_{2}}$ and subsets $\mathcal{A}_{1}\subset[n_{1}],~{}\mathcal{A}_{2}\subset[n_{2}]$, let $A_{\mathcal{A}_{1},\mathcal{A}_{2}}=\left(A_{i,j}\right)_{i\in\mathcal{A}_{1},j\in\mathcal{A}_{2}}\in\mathbb{R}^{|\mathcal{A}_{1}|\times|\mathcal{A}_{2}|}$ be a submatrix of $A$. For $j\in[n_{2}]$, define $A_{\cdot,j}=A_{[n_{1}],\{j\}}$ and $A_{\cdot,-j}=A_{[n_{1}],\{j\}^{c}}$. Define $|A|_{\infty}=\max_{i\in[n_{1}],j\in[n_{2}]}|A_{i,j}|$ to be the entry-wise maximum norm, and $\|A\|_{l}=\sup_{x\in\mathbb{R}^{n_{2}}\setminus\{0_{n_{2}}\}}\|Ax\|_{l}/\|x\|_{l}$ to be the matrix $l$-norm with $l\in[1,\infty]$. Denote by $S_{+}^{n}$ the set of $n\times n$ symmetric positive definite matrices.
@@ 2.2 Problem Setup
Our goal is to develop distributed integrative estimation procedures for the precision matrices within the space (1), while ensuring communication efficiency.
It is worth highlighting that, $s_{2}=0$ corresponds to complete homogeneity $\Omega^{(1)}=\cdots=\Omega^{(M)}$, while $s_{1}=0$ indicates the group sparsity (Yuan and Lin,, 2006; Lounici et al.,, 2011) that has been widely considered in joint precision matrix estimations (e.g., Danaher et al.,, 2014; Cai et al., 2016a, ; Ren et al.,, 2019). Therefore, the parameter space (1) is flexible and can effectively model the similarity and heterogeneity. However, this flexibility, combined with the computational and statistical challenges inherent in precision matrix estimation, further amplifies the difficulty of distributed integrative analysis.
@@ 3 Methodology
    Variance estimation: compute $\widehat{v}_{j,k}^{(m)}$ via (9) for $j,k\in[p]$.
@@
    Transmit the quantities $n_{m}$, $\widebar{\Omega}^{(m)}$ and $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ to the central site.
@@
    $\widetilde{\Omega}^{(m)}=\widehat{\Gamma}+\widehat{\Lambda}^{(m)}=(\widehat{\Gamma}_{j,k})_{j,k\in[p]}+(\widehat{\Lambda}_{j,k}^{(m)})_{j,k\in[p]},~{}m\in[M],$
    with $\widehat{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k}),~{}\widehat{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ and $\widebar{\Omega}_{j,k}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}_{j,k}^{(m)}$, where $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$, $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ are thresholding functions, and $\lambda_{1,j,k},\lambda_{2,j,k}\geq 0$ are shrinkage levels.
@@ 3.1 Individual Local Estimation
Due to the biases induced by the regularization in (3), debiased methods have been developed to make inference on the entries of $\Omega^{(m)}$ (e.g., Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018). In general, the debiased estimator $\widebar{\Omega}^{(m)}$ can be represented as follows:
@@
where $\widehat{\Sigma}^{(m)}={n_{m}^{-1}}\sum_{i=1}^{n_{m}}(X_{i}^{(m)}-\widebar{X}^{(m)})(X_{i}^{(m)}-\widebar{X}^{(m)})^{\scriptscriptstyle\sf T}$ is the sample covariance. It can be shown that $\widebar{\Omega}_{j,k}^{(m)},j,k\in[p]$ is asymptotically linear with influence functions
@@
To make valid inference for $\Omega_{j,k}^{(m)}$, the variance term $v_{j,k}^{(m)}$ should be estimated consistently. In the special case where the data distribution is jointly Gaussian, elemental calculation yields that $v_{j,k}^{(m)}=(\Omega_{j,k}^{(m)})^{2}+\Omega_{j,j}^{(m)}\Omega_{k,k}^{(m)}$, and hence plug-in estimator using either $\widehat{\Omega}^{(m)}$ or $\widebar{\Omega}^{(m)}$ can be directly employed to estimate $v_{j,k}^{(m)}$. In this article, to accommodate more general distributions beyond Gaussianity, we estimate $v_{j,k}^{(m)}$ through the empirical variance of the estimated influence functions:
@@
Finally, the $m$-th site transmits the sample size $n_{m}$, the local estimator $\widebar{\Omega}^{(m)}$ in (6), and the variance estimator $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ in (9) to the central site, followed by the integration step in Section 3.2.
@@ 3.2 Heterogeneity-Adjusted Aggregating and Thresholding
To exploit the data similarity and improve the estimation accuracy, we next perform a heterogeneity-adjusted aggregating and thresholding step based on the local estimation in the previous section.
@@
$\widetilde{\Omega}^{(m)}=\widehat{\Gamma}+\widehat{\Lambda}^{(m)},m\in[M],$ (12)
where the $\widehat{\Gamma}=(\widehat{\Gamma}_{j,k})_{j,k\in[p]}$ and $\widehat{\Lambda}^{(m)}=(\widehat{\Lambda}_{j,k}^{(m)})_{j,k\in[p]}$ are given by
$\widehat{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k}),$ (13)
$\widehat{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ (14)
@@
[…]
For the shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in (13)-(14), both universal shrinkage (e.g., Bickel and Levina,, 2008; Rothman et al.,, 2009; Cai and Zhou,, 2012) and entry-dependent adaptive shrinkage (e.g., Cai and Liu,, 2011; Fan et al.,, 2013) can be applied. In this article, we employ the entry-dependent levels to accommodate the entry-wise volatility heteroscedasticity for $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$, as shown in Propositions 1-2 of Section 4.1.
@@
The HEAT method achieves both communication efficiency in distributed settings and computation efficiency for joint precision matrix estimation. On the one hand, the dimensions of the transmitted summaries in Algorithm 3 are proportional to the dimension of the target parameters. It aligns with the existing communication-efficient algorithms developed under homogeneity (e.g., Lee et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019) and shows advantages over the heterogeneous approach in Cai et al., (2022), where $p\times p$ matrices are transmitted even though the target parameters are $p$-dimensional. On the other hand, unlike existing joint precision matrix estimation algorithms that involve highly non-smooth optimizations and computationally intensive procedures, HEAT is optimization-free for integration, and therefore is computationally efficient.

### v2

# Optimal Integrative Estimation for Distributed Precision Matrices with Heterogeneity Adjustment

@@ 1 Introduction
Nevertheless, synthesizing information from multiple sources is highly challenging. First, the prospective between-study heterogeneity complicates data integration. Such heterogeneity may manifest in various forms, including variations in genotype-phenotype associations (Frazier-Wood et al.,, 2013) and the differences in genetic variants across diverse ethnic populations (Gurdasani et al.,, 2019). Additionally, variability arising from inconsistent data collection and measurement across institutions further contributes to data heterogeneity (Kohane et al.,, 2021). Second, due to privacy (Arellano et al.,, 2018; Kohane et al.,, 2021) or communication concerns, individual raw data often cannot be shared across different locations. Instead, only summary or intermediate statistics can be communicated among different sources (Duan et al.,, 2020; Zhou et al.,, 2022). This constraint on data sharing may negatively impact the learning accuracy, exacerbating the challenges of data integration.
@@ 1.1 Related Works
In recent decades, significant progress has been made in the field of high-dimensional precision matrix estimation under sparsity, including penalized maximum likelihood procedures (Yuan and Lin,, 2007; Rothman et al.,, 2008; Lam and Fan,, 2009; Ravikumar et al.,, 2011), penalized linear regression approaches (Meinshausen and Bühlmann,, 2006; Sun and Zhang,, 2013; Ren et al.,, 2015; Liu and Luo,, 2015), and regularized estimating equation methods (Yuan,, 2010; Cai et al.,, 2011; Cai et al., 2016b, ). The non-negligible bias induced by the regularization in these methods has prompted numerous debiasing strategies aimed at statistical inference on the entries of precision matrices (Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018; Chang et al.,, 2018).
@@
In a similar context to joint estimation, there is a substantial body of literature focusing on distributed analysis under distribution homogeneity. Examples include the averaging methods (Zhang et al.,, 2015; Lee et al.,, 2017; Battey et al.,, 2018; Fan et al.,, 2019; Dobriban and Sheng,, 2020, 2021; Lv and Lian,, 2022; Tu et al., 2024b, ), and the surrogate loss function methods (Wang et al.,, 2017; Jordan et al.,, 2019) with applications in distributed optimization (Fan et al.,, 2023), simultaneous inference (Yu et al.,, 2022), non-smooth analysis (Tan et al.,, 2022) and semi-supervised estimation (Tu et al., 2024a, ). These methods achieve communication-efficiency in the sense that the dimensions of the transmitted statistics are proportional to those of target parameters. Besides, averaging aggregations under communication constraints are developed for estimation (Garg et al.,, 2014; Braverman et al.,, 2016; Zhu and Lafferty,, 2018; Szabó and van Zanten,, 2020; Cai and Wei,, 2022; Acharya et al.,, 2023, 2024) and testing (Szabó et al.,, 2022, 2023). In addition, Chen et al., (2019) smooths the non-smooth objective and solves it by transmitting statistics with dimensions that are quadratic in the parameter dimension; see also Wang et al., (2019); Chen et al., (2024) for further applications. On the other hand, when model distributions are heterogeneous while target parameters remain homogeneous across sites, distributed analysis can be achieved by averaging (Zhao et al.,, 2016; Gu and Chen,, 2023; Chen et al.,, 2025) or surrogate estimating equations (Duan et al.,, 2022).
@@ 1.2 Contributions
The newly established methodologies and theories in this article make several valuable contributions to both distributed statistical learning and integrative precision matrix analysis. First, the proposed HEAT and IteHEAT methods introduce novel methodological advances in distributed algorithms under heterogeneity. In comparison to existing communication-efficient algorithms developed under homogeneity, our approaches demonstrate flexibility and applicability across distributional heterogeneity scenarios. Additionally, our methods provide communication advantages over distributed algorithms designed for heterogeneous settings, such as Cai et al., (2022). Second, the proposed IteHEAT algorithm provides insights into the statistical refinement of higher-order errors and demonstrates potential for improving estimation accuracy in integrative analysis. Third, the proposed methods complement the field of multi-task learning for precision matrices. Unlike the typical joint estimation approaches that rely on solving highly non-smooth regularized optimizations, HEAT and IteHEAT are optimization-free for integration, ensuring computational efficiency. Fourth, we establish the statistical optimality theories of HEAT and IteHEAT by deriving the rates of convergence and the corresponding minimax lower bounds. To the best of our knowledge, the current minimax theories are the first among the extensive literature on joint precision matrix estimation. The introduced lower bound techniques and constructions of least favorable matrices are of independent interest and make valuable contributions to the field.
@@ 2.1 Notations
For an integer $k>0$, let $k]=\{1,2,\ldots,k\}$; denote by $0_{k}$ and $1_{k}$ the vectors of zeros and ones with length $k$, respectively. For a set $\mathcal{A}$, define $|\mathcal{A}|$ to be its cardinality. For a vector $a\in\mathbb{R}^{n}$ and a subset $\mathcal{A}\subset[n]$, denote by $a_{\mathcal{A}}=\left(a_{i}\right)_{i\in\mathcal{A}}\in\mathbb{R}^{|\mathcal{A}|}$ a subvector of $a$. Let $\mathcal{A}^{c}=[n]\setminus\mathcal{A}$, and define $a_{-\mathcal{A}}=a_{\mathcal{A}^{c}}\in\mathbb{R}^{n-|\mathcal{A}|}$, and $a_{-j}=a_{-\{j\}}$. For a matrix $A=\left(A_{i,j}\right)_{i\in[n_{1}],j\in[n_{2}]}\in\mathbb{R}^{n_{1}\times n_{2}}$ and subsets $\mathcal{A}_{1}\subset[n_{1}],~\mathcal{A}_{2}\subset[n_{2}]$, let $A_{\mathcal{A}_{1},\mathcal{A}_{2}}=\left(A_{i,j}\right)_{i\in\mathcal{A}_{1},j\in\mathcal{A}_{2}}\in\mathbb{R}^{|\mathcal{A}_{1}|\times|\mathcal{A}_{2}|}$ be a submatrix of $A$. For $j\in[n_{2}]$, define $A_{\cdot,j}=A_{[n_{1}],\{j\}}$ and $A_{\cdot,-j}=A_{[n_{1}],\{j\}^{c}}$. Define $|A|_{\infty}=\max_{i\in[n_{1}],j\in[n_{2}]}|A_{i,j}|$ to be the entry-wise maximum norm, and $\|A\|_{l}=\sup_{x\in\mathbb{R}^{n_{2}}\setminus\{0_{n_{2}}\}}\|Ax\|_{l}/\|x\|_{l}$ to be the matrix $l$-norm with $l\in[1,\infty]$. Denote by $S_{+}^{n}$ the set of $n\times n$ symmetric positive definite matrices.
@@ 2.2 Problem Setup
Our goal is to develop distributed integrative estimation procedures for the precision matrices within the space (1), while ensuring communication efficiency and statistical guarantees.
It is worth highlighting that, $s_{2}=0$ corresponds to complete homogeneity $\Omega^{(1)}=\cdots=\Omega^{(M)}$, while $s_{1}=0$ indicates the group sparsity (Yuan and Lin,, 2006; Lounici et al.,, 2011) that has been widely considered in joint precision matrix estimations (Danaher et al.,, 2014; Cai et al., 2016a, ; Ren et al.,, 2019). Therefore, the parameter space (1) is flexible and can effectively model the similarity and heterogeneity. However, this flexibility, combined with the computational and statistical challenges inherent in precision matrix estimation, further amplifies the difficulty of distributed integrative analysis.
@@ 1 Distributed Estimation for Heterogeneous Precision Matrices
    Transmit the quantities $n_{m}$ and $\widebar{\Omega}^{(m)}$ to the central site.
@@
    (Optional) Variance estimation: compute $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ via (9), and transmit $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ to the central site.
@@
    $\widetilde{\Omega}^{(m)}=\widetilde{\Gamma}+\widetilde{\Lambda}^{(m)}=(\widetilde{\Gamma}_{j,k})_{j,k\in[p]}+(\widetilde{\Lambda}_{j,k}^{(m)})_{j,k\in[p]},~m\in[M],$
    with $\widetilde{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k})$, $\widetilde{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ and $\widebar{\Omega}_{j,k}=\frac{1}{N}\sum_{m=1}^{M}n_{m}\widebar{\Omega}_{j,k}^{(m)}$, where $T_{1,\lambda}:\mathbb{R}\rightarrow\mathbb{R}$, $T_{2,\lambda}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{M}$ are thresholding functions, and $\lambda_{1,j,k},\lambda_{2,j,k}\geq 0$ are shrinkage levels.
@@ 3.1 Individual Local Estimation
Due to the biases induced by the regularization in (3), debiased methods have been developed to obtain asymptotically unbiased estimators (e.g., Liu,, 2013; Janková and van de Geer,, 2015, 2017; Ning and Liu,, 2017; Neykov et al.,, 2018). In general, the debiased estimator $\widebar{\Omega}^{(m)}$ can be represented as follows:
@@
where $\widehat{\Sigma}^{(m)}={n_{m}^{-1}}\sum_{i=1}^{n_{m}}(X_{i}^{(m)}-\widebar{X}^{(m)})(X_{i}^{(m)}-\widebar{X}^{(m)})^{\scriptscriptstyle\sf T}={n_{m}^{-1}}(\widecheck{X}^{(m)})^{\scriptscriptstyle\sf T}\widecheck{X}^{(m)}$ is the sample covariance. It can be shown that $\widebar{\Omega}_{j,k}^{(m)},j,k\in[p]$ is asymptotically linear with influence functions
@@
To conduct statistical inference on an entry $\Omega_{j,k}^{(m)}$, typically the variance term $v_{j,k}^{(m)}$ should be estimated consistently to quantify the uncertainty of the derived estimator $\widebar{\Omega}_{j,k}^{(m)}$. In the special case where the data distribution is jointly Gaussian, elemental calculation yields that $v_{j,k}^{(m)}=(\Omega_{j,k}^{(m)})^{2}+\Omega_{j,j}^{(m)}\Omega_{k,k}^{(m)}$, and hence a plug-in estimator using either $\widehat{\Omega}^{(m)}$ or $\widebar{\Omega}^{(m)}$ can be directly employed to estimate $v_{j,k}^{(m)}$. In this article, to accommodate more general distributions beyond Gaussianity, we propose to estimate $v_{j,k}^{(m)}$ through the empirical variance of the estimated influence functions:
@@
Finally, the $m$-th site transmits the sample size $n_{m}$ and the local estimator $\widebar{\Omega}^{(m)}$ in (6) to the central site, followed by the integration step in subsequent Section 3.2. Depending on the specific implementation of the integrative procedure, the local site may also optionally transmit the variance estimation $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ in (9). Notably, such sharing of summary statistics in communication precludes the transfer of individual-level data, and therefore provides raw data protection in multi-source data integration.
Remark 1.
The constructions of debiased estimator in (6) and variance estimator in (9) ensure that both $\widebar{\Omega}^{(m)}$ and $(\widehat{v}_{j,k}^{(m)})_{j,k\in[p]}$ are symmetric, i.e., $\widebar{\Omega}_{j,k}^{(m)}=\widebar{\Omega}_{k,j}^{(m)}$ and $\widehat{v}_{j,k}^{(m)}=\widehat{v}_{k,j}^{(m)}$ for all $j,k\in[p]$. Therefore, in practice, it is sufficient for local sites to transmit only the upper triangular entries corresponding to the index set $\{(j,k):1\leq j\leq k\leq p\}$, which reduces the communication burden.
Remark 2.
The proposed transmission achieves communication efficiency in distributed settings, as the size of transmitted summary statistics in Algorithm 3 is proportional to the dimension of the target parameters. It aligns with the existing communication-efficient algorithms designed under homogeneity (Lee et al.,, 2017; Wang et al.,, 2017; Battey et al.,, 2018; Jordan et al.,, 2019), and shows advantage over the heterogeneous approaches (Cai et al.,, 2022; Li et al.,, 2023) that require the transmission of $p\times p$ second-order Hessian matrices even when the target parameter is only $p$-dimensional. See Section LABEL:sec:more_discuss_commu of the supplement for more detailed discussions.
@@ 3.2 Heterogeneity-Adjusted Aggregating and Thresholding
To exploit the data similarity and improve the estimation accuracy, we next perform a HEterogeneity-adjusted Aggregating and Thresholding (HEAT) step based on the local estimation in the previous section to obtain integrative estimators.
@@
$\widetilde{\Omega}^{(m)}=\widetilde{\Gamma}+\widetilde{\Lambda}^{(m)},m\in[M],$ (12)
where the $\widetilde{\Gamma}=(\widetilde{\Gamma}_{j,k})_{j,k\in[p]}$ and $\widetilde{\Lambda}^{(m)}=(\widetilde{\Lambda}_{j,k}^{(m)})_{j,k\in[p]}$ are given by
$\widetilde{\Gamma}_{j,k}=T_{1,\lambda_{1,j,k}}(\widebar{\Omega}_{j,k}),$ (13)
$\widetilde{\Lambda}_{j,k}^{(\cdot)}=T_{2,\lambda_{2,j,k}}(\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}),$ (14)
@@
where $\|\cdot\|$ represents some norm for the space $\mathbb{R}^{M}$ and will be specified later in Assumption 4 of Section 4. Proper functions for $T_{2,\lambda}$ include the multivariate versions of soft, hard, SCAD and MCP thresholding; see Section LABEL:sec:addition_choice_threshold of the supplement for detailed examples.
The shrinkage levels $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ in (13)-(14) can be determined using either universal shrinkage (e.g., Bickel and Levina,, 2008; Rothman et al.,, 2009; Cai and Zhou,, 2012) or entry-dependent adaptive shrinkage (e.g., Cai and Liu,, 2011; Fan et al.,, 2013). Specifically, for universal shrinkage, the thresholds $\lambda_{1,j,k}$ and $\lambda_{2,j,k}$ remain the same across all entries $(j,k)$, while adaptive shrinkage sets distinct thresholds to account for the entry-wise volatility heteroscedasticity of $\widebar{\Omega}_{j,k}$ and $\widebar{\Omega}_{j,k}^{(\cdot)}-\widebar{\Omega}_{j,k}1_{M}$ for $j,k\in[p]$. The choice between universal and adaptive shrinkage depends on whether the local variance estimators from Section 3.1 are transmitted to the central site.
@@
The HEAT approach achieves computational efficiency for joint precision matrix estimation. As reviewed in Section 1, existing joint estimation algorithms mainly involve highly non-smooth and complex regularizations, typically leading to computationally intensive optimization procedures. In contrast, HEAT in step (12) is optimization-free for integration, which therefore renders the procedure computationally efficient at the central site.